# The Global Glimm Property for C*-algebras of topological dimension zero

### v1

@@ 1. Introduction
The Global Glimm Property always implies nowhere scatteredness, and the Global Glimm Problem asks if the converse is also true. This is known to hold under the additional assumption of real rank zero [ER06], or stable rank one [APRT22], or Hausdorff finite dimensional primitive ideal space [BK04], or finite decomposition rank ([EV25, Theorem 2.3]).
Recently, a general framework to approach this problem was given in [TV23] employing Cuntz semigroup techniques. Making use of the tools developed in [TV23], we generalize the real rank zero result from [ER06]:
[…]
@@
A $C^{*}$-algebra is said to have topological dimension zero if its primitive ideal space has a basis consisting of compact, open sets. This notion was introduced by Brown and Pedersen [BP09, Remark 2.5(vi)] as a generalization of real rank zero. More generally, it is known that every $C^{*}$-algebra with real rank zero has the ideal property (that is, its projections separate its closed ideals), and that every $C^{*}$-algebra with the ideal property has topological dimension zero. All three notions agree for separable $\mathcal{O}_{2}$-stable algebras [PR07, Corollary 4.3].
As a corollary of Section 1 we obtain that nowhere scattered $C^{*}$-algebras with finite nuclear dimension and topological dimension zero are pure (Section 2). Further, using results by Zhang [Zha90] and the third-named author [Vil23], we prove that $\sigma$-unital, purely infinite $C^{*}$-algebras of real rank zero have purely infinite multiplier algebras (Section 2).
@@ 2. The proofs
By [RT17, Lemma 7.12], for every $\varepsilon^{\prime}>0$ the set of elements $c\in A_{+}$ such that $(c-\varepsilon^{\prime})_{+}$ is pseudocompact is dense in $A_{+}$, where an element $d\in A_{+}$ is said to be pseudocompact if there exist $\delta>0$ and $n\in{\mathbb{N}}$ such that $[d]\leq n[(d-\delta)_{+}]$ in $\operatorname{Cu}(A)$. Applying this, we obtain $c\in A_{+}$ such that
@@
Pick $\delta>0$ and $n\in{\mathbb{N}}$ such that $[d]\leq n[(d-\delta)_{+}]$. Using that $x^{\prime}\ll[(a-\varepsilon)_{+}]\leq[d]$, we may assume that $\delta$ is so small that $x^{\prime}\ll[(d-\delta)_{+}]$. Set $y:=[(d-\tfrac{\delta}{2})_{+}]$ and $y^{\prime}:=[(d-\delta)_{+}]$. Then
@@
$c^{\prime}\ll e^{\prime}\ll e\ll c,\,\,\,\text{ and }\,\,\,e\ll\infty e^{\prime}.$
@@
The forward implication holds in general. For the converse, assume that $A$ is nowhere scattered. Then $\operatorname{Cu}(A)$ is weakly $(2,\omega)$-divisible by [TV24, Theorem 8.9]. Further, the Cuntz semigroup of every $C^{*}$-algebra satisfies (O5)-(O8). Since $A$ has topological dimension zero, $\operatorname{Cu}(A)\otimes\{0,\infty\}$ is algebraic by Section 2.
We can therefore apply Lemma 2.1 to deduce that $\operatorname{Cu}(A)$ is $(2,\omega)$-divisible. Then, by [TV23, Theorem 3.6], $A$ has the Global Glimm Property. ∎
@@
In [KR02, Proposition 4.15], Kirchberg and Rørdam show that a $C^{*}$-algebra is purely infinite if and only if it is weakly purely infinite and has the Global Glimm Property, and they ask if every weakly purely infinite $C^{*}$-algebra is purely infinite ([KR02, Question 9.5]). Using Section 2, we obtain a positive answer to their question for $C^{*}$-algebras with topological dimension zero. This includes $C^{*}$-algebras that are simple or have real rank zero, for which it was already shown in [KR02] that weak pure infiniteness implies pure infiniteness.
[…]

### v2

@@ 1. Introduction
The Global Glimm Property always implies nowhere scatteredness, and the Global Glimm Problem asks if the converse is also true. This is known to hold under the additional assumption of real rank zero [ER06], or stable rank one [APRT22], or Hausdorff finite-dimensional primitive ideal space [BK04], or finite decomposition rank ([EV25, Theorem 2.3]).
Recently, a general framework to approach this problem was given in [TV23] employing Cuntz semigroup techniques. Making use of the tools developed in [TV23], we solve the Global Glimm Problem for a large class of $C^{*}$-algebras:
Theorem A (2).
@@
A $C^{*}$-algebra is said to have topological dimension zero if its primitive ideal space has a basis consisting of compact, open sets. This notion was introduced by Brown and Pedersen [BP09, Remark 2.5(vi)] as a generalization of real rank zero. More generally, it is known that every $C^{*}$-algebra with real rank zero has the ideal property (that is, its projections separate its closed ideals), and that every $C^{*}$-algebra with the ideal property has topological dimension zero.
Section 1 both recovers and extends previous solutions of the Global Glimm Problem. First, since real rank zero implies topological dimension zero, our result subsumes the real rank zero case established in [ER06]. Second, because Hausdorff zero-dimensional primitive ideal space implies topological dimension zero, we partially recover the Hausdorff finite-dimensional primitive ideal space case treated in [BK04]. Beyond these cases, Section 1 solves the Global Glimm Problem for classes of $C^{*}$-algebras not covered by previous results in the literature. For example, Zhang [Zha90] showed that multiplier algebras of $\sigma$-unital, real rank zero $C^{*}$-algebras have topological dimension zero, and thus fall within the scope of Section 1. However, such multiplier algebras are not encompassed by earlier results, since they need not have real rank zero, Hausdorff primitive ideal space, or stable rank one (see Section 2).
As a corollary of Section 1 we obtain that nowhere scattered $C^{*}$-algebras with finite nuclear dimension and topological dimension zero are pure (Section 2). Further, using results by Zhang [Zha90] and the third-named author [Vil23], we prove that $\sigma$-unital, purely infinite $C^{*}$-algebras of real rank zero have purely infinite multiplier algebras (Section 2), thereby partially resolving a problem from [KR02]. We also recover the result of Elliott and Rouzbehani [ER23] that a weakly purely infinite $C^{*}$-algebra with topological dimension zero is purely infinite (Section 2).
@@ 2. The proofs
By [RT17, Lemma 7.12], for every $\varepsilon^{\prime}>0$ the set of elements $c\in A_{+}$ such that $(c-\varepsilon^{\prime})_{+}$ is pseudocompact is dense in $A_{+}$, where an element $d\in A_{+}$ is said to be pseudocompact if there exist $\delta>0$ and $n\in{\mathbb{N}}$ such that $d]\leq n[(d-\delta)_{+}]$ in $\operatorname{Cu}(A)$. Applying this, we obtain $c\in A_{+}$ such that
@@
Pick $\delta>0$ and $n\in{\mathbb{N}}$ such that $d]\leq n[(d-\delta)_{+}]$. Using that $x^{\prime}\ll[(a-\varepsilon)_{+}]\leq[d]$, we may assume that $\delta$ is so small that $x^{\prime}\ll[(d-\delta)_{+}]$. Set $y:=[(d-\tfrac{\delta}{2})_{+}]$ and $y^{\prime}:=[(d-\delta)_{+}]$. Then
@@
ceec, and $^{\prime}\ll e^{\prime}\ll e\ll c$$,\text${ and $\ll\infty e^{\prime}$.
@@
The forward implication holds in general. For the converse, assume that $A$ is nowhere scattered. Then $\operatorname{Cu}(A)$ is weakly $(2,\omega)$-divisible by [TV24, Theorem 8.9]. Further, the Cuntz semigroup of every $C^{*}$-algebra satisfies (O5)-(O8). Since AA has topological dimension zero, $\operatorname{Cu}(A)\otimes\{0,\infty\}$ is algebraic by Section 2.
We can therefore apply Lemma 2.2 to deduce that $\operatorname{Cu}(A)$ is $(2,\omega)$-divisible. Then, by [TV23, Theorem 3.6], AA has the Global Glimm Property. ∎
@@
In [KR02, Proposition 4.15], Kirchberg and Rørdam show that a $C^{*}$-algebra is purely infinite if and only if it is weakly purely infinite and has the Global Glimm Property, and they asked if every weakly purely infinite $C^{*}$-algebra must in fact be purely infinite ([KR02, Question 9.5]). This question was settled affirmatively by Elliott and Rouzbehani [ER23] for $C^{*}$-algebras with topological dimension zero, a class that includes both the simple and the real rank zero cases already treated in [KR02]. We recover their result.
Corollary 2.5 (Elliott-Rouzbehani [ER23]).